OpenStudy (anonymous):

5e^{2x + 11} = 30 I know x = -4.604 I just need the steps

OpenStudy (anonymous):

\[5e^{2x + 11} = 30\] Divide each side both five. \[e^{2x + 11} = 6\] Can you continue from there or do you need more assistance?

OpenStudy (anonymous):

ln(e^{2x + 11}) = ln(6) Right?

OpenStudy (anonymous):

Yes :-). ln cancels out e so: 2x + 11 = ln(6) 2x + 11 = 1.79175946923 Are you able to solve for x?

OpenStudy (anonymous):

2x = 1.79175946923 - 11 2x = -9.20824053077

OpenStudy (anonymous):

Divide both sides by 2: x = -4.604 right?

OpenStudy (anonymous):

Correct :-).

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :-).